OpenStudy (anonymous):

a laundry bag contains 4 green socks, 6 red socks,and 2 blue socks. Inez reaches into the bag and randomly pull out 2 socks, one right after the other. what is the probability that she pulls out a green sock and then a blue sock?

OpenStudy (amistre64):

12 socks total.... theres a 4 in 12 chances of picking a green sock.... or is that 1 in 3?

OpenStudy (ashwinram):

4 in 12 chance of green sock for the first pick

OpenStudy (ashwinram):

if you get green, now 2 in 11 chance of blue sock for second pick

OpenStudy (amistre64):

2 in 11 chance for the next one being blue

OpenStudy (anonymous):

in my multiple hoisce these are the chouces

OpenStudy (amistre64):

its like reading a stroke :)

OpenStudy (anonymous):

a. 2/11 b.6/11 c.2/33 d.8/33

OpenStudy (anonymous):

* in my multiple choices theses are the choices

OpenStudy (ashwinram):

first action: 4/12 AND second action: 2/11 so combined probability 4/12 * 2/11

OpenStudy (ashwinram):

AND means multiply for probabilities

OpenStudy (anonymous):

but how do i multiply it

OpenStudy (ashwinram):

do you know how to multiple fractions?

OpenStudy (anonymous):

no not really

OpenStudy (ashwinram):

multiply both numerators with each other

OpenStudy (ashwinram):

and multiply both denominators with each other

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 8/132

OpenStudy (ashwinram):

\[(4/12) * (3/11) = (4*3) / (12*11)\]

OpenStudy (ashwinram):

you can simplify before or after multiplying

OpenStudy (anonymous):

i got it thanks i got 2/33

OpenStudy (ashwinram):

before multiplying: simplify 4/12 which gives you 1/3. Now multiply 1/3 * 2/11 which is (1*2) / (3*11) = 2/33

OpenStudy (anonymous):

thanks for your help

OpenStudy (ashwinram):

or after multiplying: first multiply and get (4*2) / (12*11) = 8/132 and now simplify to 2/33

OpenStudy (ashwinram):

it's usually easier to simplify first

OpenStudy (ashwinram):

you're welcome :) glad i could help